# CAIRe: Cultural Attribution of Images by Retrieval-Augmented Evaluation

### v1

@@ 7 Related Work
CUBE [24] assesses cultural diversity using a manually curated reference set limited to eight countries and three domains with 300k artifacts. In contrast, our framework leverages a knowledge base spanning 6M concepts, enabling broader assessments. Unlike prior work that primarily evaluates batch-level diversity and focuses only on generated images, our approach assesses cultural relevance at the individual image level and works with both generated and natural images. Finally, while most existing evaluations assign a single country label per image, we introduce a more flexible and inclusive method, allowing multiple cultural proxies to define and assess cultural relevance.
Work on multilingualism in text-to-image models overlaps with themes of cultural evaluation. Saxon and Wang [44] introduced an evaluation of cross-lingual capabilities of a slate of T2I models using image similarity over cultural universals. This line of work has faced challenges for a lack of techniques like CAIRe to perform cultural attribution [43, 45].
Binary v/s Graded Judgements in Cultural Contexts
: Prior work across education, communication, and psychology consistently shows that cultural relevance is best understood as a matter of degree rather than a binary condition. Binary judgments oversimplify the subjective and context-dependent nature of cultural fit, potentially masking ambivalence or partial connections [51]. Empirical studies demonstrate that graded scales more effectively capture this nuance. For instance, Ebe [9] showed that elementary students rated storybooks for cultural relevance using a rubric, and those receiving higher relevance ratings led to better comprehension and engagement. Similarly, in health communication, Vereen et al. [50] found that Black women rated culturally tailored visual advertisements significantly higher on cultural relevance than generic versions, using a 7-point Likert scale, though both ads conveyed identical health information. Narrative media research also supports this approach: [32] reported that Latino viewers rated culturally resonant videos as more culturally similar to themselves, which corresponded to stronger identification with the characters. Finally, classic advertising research by Whittler [55] found that participants rated advertisements featuring spokespeople of their own race as more relatable, with these higher identification ratings predicting more favorable attitudes toward the ad. Together, these studies affirm that cultural relevance judgments are inherently graded, and that ordinal scales better capture varying levels of perceived fit and resonance. For tasks like estimating the cultural relevance of images, which demand modeling subtle human perceptions across diverse cultural identities, graded scoring (such as a 1–5 relevance scale) is therefore far more appropriate and informative than binary labels.
@@ 8 Conclusion
From our formulation of the visual cultural attribution task, we have produced two test sets that capture important desiderata: specific accuracy over rare entities and faithful modeling of human opinions over diverse cultures.
Our CAIRe family of visual cultural attribution metrics is the first framework in this space that permits assessments over free-text culture labels (open vocabulary) and provides numerical scores rather than binary judgments. By leveraging a massively multilingual, multimodal knowledge base and state-of-the-art vision-language encoders, CAIRe can account for rare entities, and using those retrieved KB entries as input to LMs for final judgment, we robustly provide these numerical scores. Our results demonstrate that CAIRe beats naive LM-based approaches on both test sets, and satisfies our need for a robust visual cultural attribution tool.
With CAIRe, we significantly expand the scope of concepts we can evaluate for cultural diversity, since we use the KB images as our reference set for grounding. Given its modularity and leverage of state-of-the-art LMs, CAIRe can continue to evolve going forward and enable research into cultural issues in multimodality.
@@ 9 Limitations & Ethical Considerations
10 Contributions
We list author contributions loosely following the CRediT author statement. [^10] https://www.elsevier.com/researcher/author/policies-and-guidelines/credit-author-statement
AY: Methodology (co-lead), Software (co-lead), Resource Development, Writing.
SY: Methodology (co-lead), Software (co-lead), Resource Development, Writing.
SK: Conceptualization (co-lead), Methodology (co-lead), Software, Validation and Analysis, Resources, Data Curation (co-lead), Writing (co-lead), Supervision (co-lead)
MS: Conceptualization (co-lead), Data Curation (co-lead), Acronymography, Methodology, Writing, Resources
GN: Supervision (co-lead), Funding Acquisition (lead), Writing.
11 Acknowledgments
We are especially grateful to Shaily Bhatt for her valuable contributions during the early stages of brainstorming for this project. We also thank the following individuals for their insightful feedback on drafts of this paper, as well as their comments during talks and presentations: members of NeuLab, Akhila Yerukola, Fernando Diaz, Danish Pruthi, Abhipsa Basu, Nithish Kannen, Partha Talukdar, and Manish Gupta.
This work was supported in part by grants from Google, the Defense Science and Technology Agency Singapore, the National Science Foundation Graduate Research Fellowship Grant No. 1650114, and the UCSB Center for Responsible Machine Learning.
@@ B.1 Alternative Cultural relevance scoring techniques
In this approach, we opt to directly use the LM’s token log-likelihoods corresponding to each candidate culture label. Given an input prompt consisting of the retrieved textual information about an entity in the image, we prompt the LM with a completion in the form of: “This text is relevant to [culture label]”. We estimate the relevance of each culture label directly using the LM head log-likelihood of the completion. In the case of VLMs, we also provide the input image as additional conditioning context in calculating the completions’ log likelihoods.
@@
^$\hat{\mathcal{L}}(c_{i},(D,I,p))=\mathcal{L}(c_{i},(D,I,p))-\lambda\cdot\max(\mathcal{L}(c_{i},(\emptyset,p)),T)$ (6)
@@ Appendix C Experimental Details
C.1 Culturally Similar Matches
@@
C.2 GlobalRG Concept List
@@
Model Image Wikipedia F1 Score
CAIRe (1-5 Scores)
Llama-3.2-11B-Vision-Instruct 48.8
Qwen2.5-7B-Instruct 77.7
Qwen2.5-VL-7B-Instruct 77.7
Pangea-7B-hf 46.6
CAIRe (Log-probabilities)
Llama-3.2-11B-Vision 65.7
Qwen2.5-7B 58.6
Qwen2.5-VL-7B-Instruct 45.1
Pangea-7B-hf 62.6
Table 10: F1-scores with Gold context
Model Image Ratio
CAIRe (1-5 Scores)
Llama-3.2-11B-Vision-Instruct 97.1
Qwen2.5-7B-Instruct 88.7
Qwen2.5-VL-7B-Instruct 84.3
Pangea-7B-hf 92.1
CAIRe (Log-probabilities)
Llama-3.2-11B-Vision 94.1
Qwen2.5-7B 89.6
Qwen2.5-VL-7B-Instruct 96.7
Pangea-7B-hf 89.5
Table 11: Ratio between CAIRe and Gold context
C.3 Cultural Relevance Scoring Criteria
@@
C.4 Example Prompt for 1-5 Scoring
@@
C.5 Upper Bound Analysis
We obtain the results in Table 10 by replacing the retrieved Wikipedia pages with the ground-truth pages corresponding to the entities in our specific set. By comparing the F1-scores of our retrieved pages to those of the gold-standard pages, we quantify the retrieval-induced error in our pipeline. Specifically, we compute the ratio of the F1-scores from our VEL system to those obtained using the gold-standard pages. As shown in Table 11, our method consistently achieves scores comparable to the gold standard, highlighting the effectiveness of our retrieval system in identifying culturally relevant entities.

### v2

@@ 7 Related Work
CUBE [23] assesses cultural diversity using a manually curated reference set limited to eight countries and three domains with 300k artifacts. In contrast, our framework leverages a knowledge base spanning 6M concepts, enabling broader assessments. Unlike prior work that primarily evaluates batch-level diversity and focuses only on generated images, our approach assesses cultural relevance at the individual image level and works with both generated and natural images. Finally, while most existing evaluations assign a single country label per image, we introduce a more flexible and inclusive method, allowing multiple cultural proxies to define and assess cultural relevance. Work on multilingualism in text-to-image models overlaps with themes of cultural evaluation. Saxon and Wang [42] introduced an evaluation of cross-lingual capabilities of a slate of T2I models using image similarity over cultural universals. This line of work has faced challenges for a lack of techniques like CAIRe to perform cultural attribution [41, 43].
@@ 8 Conclusion
In this work we introduce CAIRe, a novel and generalizable framework for visual cultural attribution. CAIRe evaluates an image’s relevance to each user-defined culture independently, rather than assigning a single, binary country label to an image as in prior work. CAIRe first uses strong VLMs to identify culturally salient entities represented in the image, grounding it to a large-scale multimodal knowledge base. Using this retrieved information about the image, CAIRe assigns final cultural relevance scores to an image for each culture label, on a scale of 1-5. CAIRe is flexible to a wide range of cultural definitions and applicable across both natural and generated images.
Our experiments show that CAIRe performs strongly in practice: it correlates well with human judgments, outperforms existing baselines across both specific and universal concept sets, and remains robust across diverse cultures. Beyond per-image attribution, we demonstrate that CAIRe functions effectively as a batch-level diversity metric, revealing cultural representation patterns in text-to-image model outputs. Together, these results establish CAIRe as a practical and scalable tool for assessing cultural relevance in vision–language systems.
@@ B.1 Alternative Cultural relevance scoring techniques
In this approach, we opt to directly use the LM’s token log-likelihoods corresponding to each candidate culture label. Given an input prompt consisting of the retrieved textual information about an entity in the image, we prompt the LM with a completion in the form of: “This text is relevant to [culture label]”. We estimate the relevance of each culture label directly using the LM head log-likelihood of the completion.
@@
$\hat{\mathcal{L}}(c_{i},(D,I,p))=\mathcal{L}(c_{i},(D,I,p))-\lambda\cdot\max(\mathcal{L}(c_{i},(\emptyset,p)),T)$ (6)
@@ Appendix C Experimental Details
C.1 Performance of VEL on the specific set
In this section, we evaluate the effectiveness of CAIRe in visual entity linking i.e., matching images to the right entities.
To estimate an upper bound on how much VEL performance impacts the overall F1 score in the case of the specific set, we incorporate gold-standard Wikipedia pages as context in our cultural relevance scoring, allowing us to quantify retrieval-induced loss.
Model Image Ratio
CAIRe (1-5 Scores)
Llama-3.2-11B-Vision-Instruct 97.1
Qwen2.5-VL-7B-Instruct 84.3
Pangea-7B-hf 92.1
CAIRe (Log-probabilities)
Llama-3.2-11B-Vision-Instruct 94.1
Qwen2.5-VL-7B-Instruct 96.7
Pangea-7B-hf 89.5
Table 9: Ratio between CAIRe and Gold context
We obtain the results in Table 10 by replacing the retrieved Wikipedia pages with the ground-truth pages corresponding to the entities in our specific set. Specifically, we compute the ratio of the F1-scores from our VEL system to those obtained using the gold-standard pages. As shown in Table 9, our method consistently achieves scores comparable to the gold standard, highlighting the effectiveness of our retrieval system in identifying culturally relevant entities.
Model Image Wikipedia F1 Score
CAIRe (1-5 Scores)
Llama-3.2-11B-Vision-Instruct 48.8
Qwen2.5-VL-7B-Instruct 77.7
Pangea-7B-hf 46.6
CAIRe (Log-probabilities)
Llama-3.2-11B-Vision 65.7
Qwen2.5-VL-7B-Instruct 45.1
Pangea-7B-hf 62.6
Table 10: F1-scores with Gold context
C.2 Culturally Similar Matches
@@
C.3 GlobalRG Concept List
@@
C.4 Cultural Relevance Scoring Criteria
@@
C.5 Example Prompt for 1-5 Scoring
@@ C.7 Alternate Thresholds
Table 14 provides the F1-score, Precision and Recall at different score thresholds for binary classification on the specific set. We note that drops in F1-score are primarily driven by drops in Precision.
@@
C.8 Alternate VL-Encoders
We run Baseline-1 using OpenCLIP variants trained on the DFN datasets [11]. Specifically, the DFN-2B variant achieves a score of 7.0% [^9] https://huggingface.co/apple/DFN2B-CLIP-ViT-L-14, the DFN-5B variant achieves 8.5% [^10] https://huggingface.co/apple/DFN5B-CLIP-ViT-H-14, and the DFN-5B-378 variant achieves 8.9% [^11] https://huggingface.co/apple/DFN5B-CLIP-ViT-H-14-378. These models outperform CLIP ViT-B/32, but achieve lower F1-scores than mSigLIP. Moreover, our approach is inherently modular and can therefore seamlessly integrate with the latest and strongest vision-language encoders as they become available.
C.9 Human Annotation Details
We initially hired 20 annotators per country. After filtering based on annotation quality and completion, the final number of usable annotators varied across countries, as summarized in Table 15 for both the universal-generated and universal-retrieved subsets. For each image, we obtain 10 relevance scores from CAIRe — one for each country as the cultural label, and 10 human scores — computed by averaging responses from annotators of that country to our relevance prompt (based on a 5-point Likert scale). These paired human and model scores are then used to compute the image-level Pearson correlation.
Country universal-generated universal-retrieved
Brazil 20 14
China 9 15
Egypt 9 13
Germany 9 15
India 20 15
Indonesia 9 14
Mexico 19 14
Nigeria 17 15
Russia 10 15
United States 20 15
Table 15: Number of annotators per country used for the two subsets (after filtering).